OpenStudy (anonymous):

How do you turn this into standard form ? 4(cos120° +i sin120° ) ?

OpenStudy (asnaseer):

use the fact that:\[\cos(120)=-\frac{1}{2}\]and:\[\sin(120)=\frac{\sqrt{3}}{2}\]